### v2

# Leveraging Diffusion Models for Parameterized Quantum Circuit Generation

Abstract
Quantum computing holds immense potential, yet its practical success depends on multiple factors, including advances in quantum circuit design. In this paper, we introduce a generative approach based on denoising diffusion models (DMs) to synthesize parameterized quantum circuits (PQCs). Extending the recent diffusion model pipeline of  [1], our model effectively conditions the synthesis process, enabling the simultaneous generation of circuit architectures and their continuous gate parameters. We demonstrate our approach in synthesizing PQCs optimized for generating high-fidelity Greenberger–Horne–Zeilinger (GHZ) states and achieving high accuracy in quantum machine learning (QML) classification tasks. Our results indicate a strong generalization across varying gate sets and scaling qubit counts, highlighting the versatility and computational efficiency of diffusion-based methods. This work illustrates the potential of generative models as a powerful tool for accelerating and optimizing the design of PQCs, supporting the development of more practical and scalable quantum applications.
Index Terms:
Quantum Circuit Synthesis, Parameterized Quantum Circuits, Generative Models, Quantum Computing, Quantum Architecture Search, Diffusion Models.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
The promise of quantum computing depends on the efficient design of quantum circuits that can be executed on currently available hardware. In the present noisy intermediate-scale quantum (NISQ) [2] era, where this challenge is most acute, circuit synthesis must account for hardware constraints such as limited qubit counts, restricted gate sets, and connectivity, while still achieving target performance metrics [3, 4, 5]. The design of an appropriate ansatz for a given task – be it approximating a target unitary, solving an optimization problem or learning from data for a quantum machine learning (QML) task – is challenging and typically relies on human expertise or brute-force search. In recent years, there has been growing interest in leveraging heuristic optimization and machine learning (ML) techniques to automate quantum circuit design. Prior approaches include evolutionary strategies [6, 7, 8, 9], reinforcement learning [10, 11, 4, 12], Bayesian optimization [13], and adaptive algorithms [14, 15]. While these methods can discover novel circuit designs, they typically require repeated evaluations of quantum metrics or cost functions during training [16][17]. This evaluation process becomes intractable for large systems and trainable parameterized quantum circuits (PQCs), creating a significant computational bottleneck for automated circuit synthesis methods [18]. Moreover, techniques trained on specific quantum tasks or circuit layouts often show poor generalization to new scenarios, as they tend to encode task-specific features rather than learning transferable principles [19]. To address these challenges, [1] introduce the idea of applying denoising diffusion models (DM) to quantum circuit synthesis. The proposed text-conditioned diffusion pipeline generates gate-based circuits for a given target objective without explicitly simulating quantum dynamics during training. While this work established DMs as a promising tool for quantum circuit synthesis, an important open challenge is handling continuous gate parameters. Commonly used rotation gates, such as single-qubit or controlled rotations, are parameterized by continuous values and significantly impact the functionality of PQCs.
In this paper, we introduce an extended DM-based pipeline that integrates continuous-valued gate parameters alongside discrete gate choices into the generation process. We train the proposed model architecture on extensive datasets of quantum circuits with varying gate sets, designed for state preparation and QML classification tasks. Experimental results demonstrate that the described methodology effectively bypasses the computationally expensive parameter training phase during the search, addressing a key limitation of many existing approaches to the the automated design of PQCs, and shows strong generalization ability to new scenarios. At inference time, the model driven by a textual prompt produces PQCs that, when subject to specified constraints, are immediately deployable on the target quantum device.
The remainder of this paper is structured as follows. Section II discusses related work. Section III presents our approach for synthesizing PQCs, detailing the diffusion-based pipeline and training procedure. Experimental results are presented in Section IV, with an analysis of potentials and limitations in Section V. Finally, Section VI concludes the paper and provides an outlook on future research.
II Related Work
The approach introduced in this work extends [1], which treat quantum circuits as structured data samples and train a generative denoising diffusion model (DM) to design new circuit designs. DMs generate synthetic data by iteratively denoising random noise samples through a learned reverse diffusion process. Introduced by [20], the fundamental principle involves adding Gaussian noise to training samples as forward Markovian diffusion process, then training a neural network to reverse noising procedure and approximate the data distribution’s score function. After training, data generation begins from a pure noise sample, and the trained model iteratively predicts and removes noise, step by step, resulting in samples resembling those from the training dataset. DMs employ denoising autoencoder architectures implemented through neural networks, typically structured to predict either the original noise-free data or the noise itself. Recent advances, such as latent diffusion and classifier-free guidance, have made DMs a leading approach in conditional generative modeling tasks across numerous domains, including image and video generation [21, 22], audio synthesis [23], and protein structure prediction [24]. Extending these advances, [1] explored the potential of DMs beyond classical domains and demonstrated their effectiveness in the context of quantum circuit synthesis. Specifically, the authors applied a denoising DM to two quantum computing tasks: (1) entanglement generation, where the model produced circuits that realize specific multi-qubit entangled states, and (2) unitary compilation, where the model generated circuits approximating random target unitaries using constrained gate sets. The DM excelled in generating valid, novel circuits and enabled advanced capabilities such as masking and editing parts of a circuit post hoc without retraining. Notably, the approach eliminates the need for classical simulations to compute cost functions comparing generated circuits to target unitaries. Ref. [25] proposes an architectural enhancement by replacing the standard U-Net used for noise prediction with a U-Net-style Diffusion Transformer. A recent concurrent work [26] applied a diffusion model to generate ansätze for variational quantum eigensolver (VQE), with a focus on quantum chemistry problems. The model was trained on a corpus of unitary coupled cluster (UCC) ansatz circuits encoded into images and successfully proposed new circuits that achieved comparable ground-state energies in VQE simulations. However, the techniques proposed in earlier studies represent gates using fixed-dimensional vectors drawn from a finite embedding set [27], limiting its ability to express parametrized quantum operations. In this work, we extend the methodology to enable the generation of PQCs with continuous gate parameters.
Beyond diffusion-based models, this work also relates more broadly to the field of quantum architecture search (QAS) research, where the goal is to automatically generate PQCs tailored to specific tasks and hardware constraints [28]. While the most approaches proposed for QAS treat the design of PQCs as a two-step process [19, 4, 17, 13], i.e., first selecting discrete gate sequences and then using classical optimizers to determine continuous parameters, several recent methods aim to learn both circuit structures and parameterizations simultaneously, optimizing over both discrete gates and continuous parameters in a unified pipeline. Ref. [29] incorporates the parameter optimization task into reinforcement learning settings enabling continuous actions. The generative quantum eigensolver (GQE) proposed in [30] is a transformer-based approach for generating of PQCs. This approach is designed for chemical problem of searching for the ground state of the electronic structure Hamiltonians of several molecules. Similarly to our approach, it integrates continuous optimizable gate parameters into a classical generative model. However, unlike the transformer-based model in GQE, we employ a denoising diffusion model for PQCs generation, which iteratively refines random noise samples to generate circuits, offering advantages in generating diverse quantum circuits with novel structures.
III Methods
In the following, we present the proposed DM-based pipeline for the generation of PQCs, detailing the circuit encoding strategy, model architecture, and training procedure.
III-A Circuit Encoding
Figure 1: Overview of our diffusion-based PQC synthesis pipeline, inspired by [1] (a) Circuit Encoding: A quantum circuit is converted into a tensor by concatenating the embeddings of discrete gates and continuous parameters (e.g., rotation angles). (b) Forward Diffusion: Gaussian noise is iteratively added to this tensor, simulating a corruption process. (c) Textual Conditioning: A pretrained CLIP encoder converts task-specific prompts into embeddings. (d) Reverse Diffusion: A U-Net with cross-attention receives the noisy tensor and text embedding, iteratively removing noise to restore gate identities and parameters. (e) Inference and Decoding: At inference, random noise is denoised under the given prompt, and the final tensor is mapped back into a readable quantum circuit.
To train our diffusion model, we represent each PQCs using two structured tensors, as depicted in Fig. 1. The first tensor is three-dimensional, where one dimension corresponds to qubit indices $N$, another to discrete time steps $T$, and the third dimension encodes discrete gate identities $d_{c}$. For each discrete times $t_{i}\in T$ step, only one gate per qubit is allowed. If no gate acts on a qubit at a particular timestep (i.e., an idle operation), we assign a special no-op token. To convert this discrete representation into a continuous embedding suitable for the diffusion model, each discrete gate type is transformed into a fixed-length embedding vector determined before training. This embedding process results in a continuous tensor representation of gates across qubits and time steps. To handle gates with continuous parameters, e.g., a rotation gate $rx(\theta)$, we introduce a second tensor with dimensions corresponding to qubit indices, discrete time steps, and the maximal number of normalized parameters used by any gate in the circuit $d_{p}$. In this parameter tensor, we encode the numerical values of gate parameters, placing a no-op token for gates without parameters. By concatenating these two tensors, each gate operation at every qubit and timestep is uniquely encoded by its discrete type embedding and its continuous parameter. As a result, identical gates with identical parameters map to identical points in this combined embedding space, whereas gates differing in type or parameter values map distinctly. This method naturally extends to gates with multiple parameters. During decoding, the combined tensor is split into a circuit matrix and a parameter matrix. For each qubit–timestep position in the circuit matrix, we use cosine similarity to match the generated gate embedding to its closest counterpart among the known gate embeddings. The corresponding parameter matrix is then applied only when a gate requires parameters, using the remaining $d_{p}$ dimensions of the vector as a parameter encoding. In principle, if a gate supports multiple parameters (not present in our current gate sets), one could introduce an inverse mapping or train an auxiliary decoder network to handle this more complex parameter space.. If the chosen gate $G$ is non-parameterized (e.g., an ’h’ gate), the parameter subvector is simply ignored.
III-B Model Architecture
In this work, we extend the quantum circuit synthesis pipeline described by [1]. Specifically, we introduce textual conditioning via a pre-trained CLIP text encoder - originally trained to align visual and textual representations through Contrastive Language-Image Pre-training (CLIP) - to guide our diffusion model towards generating PQCs optimized for specific tasks, i.e., quantum state fidelity or classification accuracy maximization. Task-specific textual prompts, for instance "Generate GHZ fidelity: $\mathsf{fid\_value}$” or "Generate Accuracy: $\mathsf{acc\_value}$”, are encoded into continuous embeddings to condition the diffusion process.
The core of our diffusion model utilizes a U-Net architecture, widely adopted in state-of-the-art diffusion-based image synthesis. The model accepts as input a noisy quantum circuit tensor of dimensions $(d_{c}+d_{p})\times N\times T$, representing embedding dimensions for gates and parameters, qubits and time steps, respectively. It outputs a noise-residual tensor of matching shape. To enable textual conditioning, we augment the standard U-Net structure with cross-attention layers, integrating CLIP-generated textual embeddings. This design explicitly conditions the generated circuits on task-specific textual prompts, eliminating the need for explicit labels of fidelity or accuracy, which the model implicitly learns from training data alone. Note that currently our approach supports parameter embeddings with $d_{p}\leq 1$.
III-C Training Strategy
Our DM is trained by simulating a forward diffusion process on quantum circuit data and learning its reverse denoising counterpart in a manner consistent with DDPM  [31] procedures. Each training iteration involves selecting a quantum circuit encoding $x_{0}$, adding Gaussian noise at a randomly sampled diffusion timestep $t$ to obtain a noised version $x_{t}$, and training the U-Net to predict this added noise, $\hat{\epsilon}_{\theta}(x_{t},t,\texttt{prompt})$. Model parameters are optimized by minimizing the mean-squared error $\mathcal{L}~{}\;=~{}\;\mathbb{E}_{x_{0},\epsilon,t}\bigl{[}\|\epsilon~{}\;-~{}\hat{\epsilon}_{\theta}(x_{t},t,\texttt{prompt})\|^{2}\bigr{]}.$
The U-Net architecture comprises an encoder-decoder convolutional network with residual blocks and transformer-based spatial attention layers. It processes a single unified tensor of size $(d_{c}+d_{p})\times N\times T$, where each slice along the embedding dimension encodes a gate identity (or no-op) and, if applicable, its continuous parameter. We incorporate the diffusion timestep $t$ via a sinusoidal positional encoding added within the residual blocks, allowing the network to learn how to invert the noise process at every step.
Conditioning on textual prompts (e.g., ‘”Generate GHZ fidelity: 1.0000”‘) is done through cross-attention layers, where the prompt is embedded by a CLIP encoder (output size $77\times 512$). For numerical stability, fidelity prompts are rounded to four decimal places, while ML accuracy prompts use one decimal place. During training, with a 10% probability, the prompt is replaced by an empty token. This “unconditional” path enables  classifier-free guidance (CFG) at inference, whereby we can smoothly interpolate between conditional and unconditional predictions  [32]. We train for $T=1000$ diffusion steps using Adam with a one-cycle learning-rate schedule.
At inference time, we draw Gaussian noise as an initial circuit encoding and iteratively denoise it using the trained U-Net. When classifier-free guidance is enabled, we selectively mix conditional and unconditional predictions to control how strictly the model follows the prompt. The final denoised tensor is decoded into a human-readable quantum circuit: we pick the closest discrete gate embedding at each position and apply the predicted continuous parameters (if any) to obtain the final PQC.
IV Results
In this section, we present the experimental results, evaluating the proposed approach and analyzing its performance across two common quantum computing tasks: quantum state preparation and a QML task focused on linear classification.
IV-A Dataset
A key element of our approach is the construction of large training datasets consisting of quantum circuits, which the diffusion model leverages to learn the circuit performance distribution. For two tasks we address, i.e., quantum state preparation and classification accuracy optimization, we build on the existing SQuASH [^1] Availibale at https://github.com/SQuASH-bench/SQuASH dataset [33]. Specifically, for the quantum state preparation, the dataset includes quantum circuits designed to generate a $3$-qubit GHZ state within a predefined approximation error. The data were collected from various search strategies used to generate circuits for the task, including random search, reinforcement learning, and an evolutionary approach. To demonstrate the model’s adaptability to different hardware constraints, we define two distinct data subsets, each characterized by a different gate alphabet. The first subset incorporates PQCs constructed using the gate set $gs1$ = {’cx’, ’h’, ’rx’, ’ry’, ’rz’, ’id’}. The second subset uses  $gs2$ = {’cz’, ’id’, ’rx’, ’rz’, ’rzz’, ’sx’, ’x’}, which corresponds to the native gate set of the IBM quantum processing unit (QPU) ’ibm_fez’.
We observed that strategically balancing the GHZ-task dataset with respect to both fidelity distribution and circuit length yielded substantial improvements in model performance. Specifically, the dataset was constructed with 25% of circuits having fidelity below 0.9, balanced evenly across different circuit lengths, while the remaining 85% consisted of circuits with high fidelity ($>0.9$) included without additional length balancing. Furthermore, we restricted the number of gates per circuit, using 3–16 gates for $gs1$ and 3–24 gates for $gs2$ and the  ML optimization task. Empirically, this targeted balancing approach notably improved training stability and enhanced the quality of generated quantum circuits.
For the machine learning setting, SQuASH provides a dedicated set of quantum circuits along with their performance on the LinearlySeparable dataset introduced in [34], which is a linear classification task. As in the previous task, we define a specific gate set, {’cx’, ’h’, ’rx’, ’ry’, ’swap’, ’crx’, ’cry’} and strategically balance the dataset in terms of classification accuracy and circuit length, which significantly improves training performance. Similar to the quantum state preparation, we constructed the dataset so that 25% of circuits had accuracy below 0.75, evenly balanced across various circuit lengths, while the remaining 75% consisted of high-accuracy circuits ($>0.75$) included without further length balancing. Empirical results confirm that this targeted dataset composition notably enhances dataset quality,
IV-B GHZ Task:
Figure 2: Comparison of fidelity distributions and evaluation metrics for two different gate sets: gs1 (top) and gs2 (middle) and generalization ability of our DM on 4 qubits for gs1 (bottom). Shown are the fidelity scores of synthesized PQCs under varying guidance scales (left) and their respective evaluation results, e.g. invalid circuits and uniqueness (right).
In this section, we present the results obtained by applying our DM to the GHZ state-preparation task under two distinct gate sets, namely $gs1$ and $gs2$. For each task, we synthesize 500 samples and 250 for 4-qubits on $gs1$, thereby ensuring statistically robust insights. As illustrated in Fig. 2, we analyze key metrics, including fidelity distributions, error counts, structural uniqueness, and parameter uniqueness. Here, parameter uniqueness refers to the diversity in continuous gate parameters, distinguishing otherwise topologically identical circuits, thereby enabling distinct quantum operations. Additionally, we discuss the critical role of  CFG in modulating circuit synthesis, leading to increasingly accurate and diverse PQCs.
For $gs1$ Fig. 2 (top) our DM consistently achieves near-perfect fidelity on three-qubit GHZ circuits, effectively converting upwards of 95% of its initial continuous tensors into valid, single-gate-per-timestep circuits. The introduction of moderate guidance scales (e.g., 10) pushes fidelity distributions toward the ideal value of $1.0$. Crucially, even at perfect fidelity targets, the model produces diverse circuit topologies. This structural variability is invaluable for practical considerations such as noise-resilience testing and hardware calibration.
An equally important observation is the model’s capacity to produce structurally diverse high-fidelity circuits. Even when tasked with the same target fidelity, the DM naturally explores multiple circuit topologies, demonstrating a talent for balancing performance with variability that could be beneficial in tasks like noise-robustness analysis or hardware calibration. Although trained solely on three-qubit examples, the DM exhibits promising scalability to four- and five-qubit scenarios Fig. 2 (bottom). At four qubits, around 30% of generated circuits reach perfect fidelity, but with a significant higher error count. At five qubits, only a few circuits achieve the ideal GHZ state, and the error count grows substantially. Furthermore, our results indicate promising scalability. The overall computational overhead remains low, enabling the rapid generation of large circuit ensembles and facilitating efficient selection of high-fidelity designs. This capability to generalize beyond the qubit counts seen in training underscores the fundamental adaptability of diffusion-based generative models. Nevertheless, while out-of-distribution fidelity declines as qubit count increases, these findings motivate further refinements—including curriculum learning over multi-qubit data or advanced sampling strategies to bridge performance gaps in higher-dimensional spaces. By systematically enhancing the training pipeline, we anticipate even more robust GHZ synthesis at larger qubit scales.
Compared to $gs1$, the $gs2$ alphabet incorporates hardware-native gates, such as ’sx’ and ’rzz’, complicating the synthesis task. As shown in Fig. 2 (middle), the absence of CFG initially yields circuits spread across lower fidelity values, reflecting the increased difficulty of orchestrating these more complex building blocks. However, once CFG is introduced, fidelity improves dramatically. At a guidance scale of 5, the DM already reaches high fidelity in mostof circuits though with 50 higher error count and less unqiues in structure. Scaling guidance to 15 pushes fidelity even further, increases the accuracy. This jump demonstrates how CFG can effectively counterbalance the tougher hardware constraints and high-dimensional parameter spaces inherent in $gs2$.
A particularly notable discovery is the emergence of layer-like circuit motifs, where single-qubit rotations precede entangling gates, illustrated in Fig. 3. The DM autonomously learns these structural templates, apparently having distilled key architectural heuristics from the training data. Importantly, each “layer” is composed of parameterized gates with unique angle settings, illustrating that diversity is not merely confined to gate ordering but also extends into the continuous parameter space.
The strong performance of our DM in synthesizing hardware-native circuits suggests a promising pathway for practical, device-level quantum compilation. By adjusting guidance scales and training on broader gate sets, one could systematically adapt the pipeline to specific backend constraints, which is an important step forward in bridging the gap between algorithmic design and near-term quantum hardware.
Figure 3: Representation of two example circuits generated by our diffusion model illustrating how the model arranges multi-layer gate structures for GHZ tasks. synthesized circuit from $gs2$ with high fidelity guidance (left) and a circuit intended for QML classification with high accuracy guidance (right).
Together, these results showcase the DM’s notable adaptability and scalability, demonstrating strong performance under stricter device constraints. Even for gate sets that demand more intricate strategies, the diffusion model augmented by CFG continues to discover high-fidelity, structurally sophisticated circuits, firmly establishing its promise for automated, hardware-aware quantum circuit synthesis.
IV-C ML task
Figure 4: Output distribution to match different accuracy targets demonstrating the models adaptability and controlled circuit generatio: Synthesized PQCs with target accuracy of 0.2 (left) and synthesized PQCs with target 0.7.
Beyond GHZ-state preparation, our diffusion model demonstrates promising potential in  QML classification tasks. To evaluate its performance, we generated 250 circuits explicitly conditioned on distinct accuracy targets (e.g., 0.7 or 0.2), assessing the model’s ability to synthesize diverse PQCs that align closely with user-defined performance levels. Notably, the dataset posed additional challenges in achieving extremely high accuracy, as it contained fewer circuits with accuracy scores of 0.9 or above. Nevertheless, the model successfully guided circuit generation toward intermediate accuracy values, effectively producing circuits centered around target accuracies such as 0.7 Fig.4 (left) and 0.2 Fig.4 (right). Leveraging CFG, our approach demonstrates the capacity to systematically explore a broad range of PQC designs, enabling precise control and rapid generation of task-oriented quantum classifiers without the need for extensive retraining or iterative simulations.
Crucially, even PQCs targeting lower accuracies maintain structurally unique designs, emphasizing the model’s extensive exploration capability. This controlled generation capability significantly advances automated PQC design, providing users a powerful tool to systematically explore circuit architectures tailored to specific quantum machine learning tasks. The diffusion model thus establishes a computationally efficient and versatile approach for discovering novel and practical quantum classifier solutions, greatly expanding the landscape of achievable quantum computational functionalities.
Taken together, these outcomes highlight a pivotal advancement in automated PQC design for both state-preparation and QML objectives. Whether aiming for high-fidelity GHZ circuits or precisely targeted classifier accuracies, the diffusion model - amplified by CFG - delivers robust, controllable, and scalable circuit synthesis. Its generalization capability and low computational overhead further solidify its practicality, as users can swiftly generate large ensembles of candidate solutions and select those that best suit their particular hardware or application constraints. Such versatility not only enriches quantum algorithm design workflows but also marks an important step toward broader adoption of generative techniques in quantum computing research and industry.
V Discussion
Our findings confirm that diffusion models can effectively synthesize high-quality PQCs without costly quantum-mechanical simulations or iterative parameter optimization. Not only does the model consistently meet or surpass the benchmarks of traditional generative methods, but it does so while maintaining considerable adaptability. As demonstrated in our experiments, CFG emerges as a crucial factor, enabling users to smoothly adjust fidelity or accuracy targets and quickly explore the resulting diverse architectural landscapes. Moreover, by simply editing textual prompts or gating configurations, the model accommodates a broad range of hardware constraints and design preferences, mirroring the flexibility of diffusion-based systems in other domains. This inherent malleability paves the way for advanced editing operations such as circuit masking or partial design overwrites, extending well beyond the usual boundaries of classical search or reinforcement learning approaches. Altogether, these attributes suggest that diffusion models not only excel in near-term tasks like GHZ-state generation or quantum classification but may also serve as robust, general-purpose frameworks for a variety of quantum synthesis challenges.
VI Conclusion and Outlook
In summary, we have introduced and evaluated a diffusion-based pipeline for generating parameterized quantum circuits across a diverse array of tasks and hardware settings. Our empirical results underscore its capacity to deliver high-fidelity GHZ states, optimize QML classifiers, and flexibly integrate hardware-related design constraints. Crucially, the method requires no iterative parameter fitting or direct quantum-state simulation during generation, offering a scalable, data-driven alternative to conventional synthesis strategies.
Looking ahead, future work could involve deeper exploration of multi-qubit scaling, integration of noise-mitigation strategies, and expanded application to broader classes of variational quantum algorithms. The apparent synergy between classifier-free guidance and diffusion modeling also opens new avenues for task-specific customizations—enabling nuanced trade-offs between circuit depth, fidelity, and hardware connectivity. We anticipate that these developments will further consolidate diffusion-based models as a cornerstone in automated quantum circuit design, driving innovation in both academic and industrial quantum research.